OpenStudy (anonymous):

logarithm help..... anyone.. how do you expand log(3v/w^2) ??

OpenStudy (phi):

Did you try?

OpenStudy (anonymous):

did i try what.. i dont know how to do it..

OpenStudy (phi):

do you remember you can change division inside a log to subtracting logs: log(a/b)= log(a)-log(b) so if you match 3v to a and w^2 to b, can you re-write the log as subtraction?

OpenStudy (anonymous):

when i write that out, is that expanded fully

OpenStudy (phi):

no, but the way to do these problems is step by step. What did you get?

OpenStudy (anonymous):

log(3v/w2)=log(3v)-log(w2)

OpenStudy (phi):

yes. Now the other rule: you can change multiplication inside a log to adding logs: log(ab)= log(a)+log(b) so, match 3 to a and v to b, and rewrite log(3v) as the sum of two logs.

OpenStudy (anonymous):

so log(3v) = log(3) + log (b)

OpenStudy (anonymous):

log(v)*

OpenStudy (phi):

yes, if you mean log(3) + log(v) (we match a to 3 and v to b to use the rule) Now -log(w^2) we can rewrite w^2 as w*w and use the same multiplication rule: -( log(w)+ log(w)) which is the same as -2log(w) (add them up, then multiply by the -1 out front) final answer: log(3)+ log(v)-2log(w)

OpenStudy (anonymous):

whoa. ok so sometimes we have to use multiple properties to get the expansion..

OpenStudy (phi):

or we could use a 3rd rule: log(a^n) (a to the nth power) = n log(a) using this rule: -log(w^2)= -2log(w) (as it should be) so those are pretty much the only rules you need to learn to do these problems

OpenStudy (phi):

If you practice, it gets easier.

OpenStudy (anonymous):

with this third rule, what do i need to do with -log(w^2)=-2log(w)

OpenStudy (phi):

the 2 in w^2 is the exponent. you move the exponent out of the log, and multiply by it. log(a^n) --> n*log(a) is is short-hand for doing this: say n=4 (as an example) a^4 = a*a*a*a (a times itself 4 times) log(a^4)= log(a*a*a*a)= log(a)+log(a)+log(a)+log(a)= 4log(a)

OpenStudy (anonymous):

ok that makes sense. ill need to practice. thanks a lot

OpenStudy (phi):

good luck!